# Hierarchical autoregressive neural networks in three-dimensional statistical system

### v1

@@ 3.1 Hierarchical autoregressive networks
The procedure for $L=16,32,\ldots$ follows the same algorithm. Going from $k$ to $k+1$ (where $L=2^{k}$), requires: i) adding (in the second stage) a new network and ii) increasing the size of the first network which fixes boundary spins (red).
The scaling with system size of the 3D hierarchical algorithm is $\sim L^{6}$: this is easy to see since the boundary and cross-like structures grow with $L$ as a surface (not a volume), hence $\sim L^{2}$. Then, we have the scaling of the HAN 3D the same as the scaling of the VAN 2D - this is even more understandable if one notices that the architecture of the boundary network (red spins in Figure 1) is just the VAN 2D. Summarizing, in 3D, the $\sim L^{9}$ scaling of the VAN is reduced to $\sim L^{6}$ of the HAN.
We used the same hyperparameters for the VAN and HAN networks: we always use dense networks with two layers, an ADAM optimizer was chosen, and the learning rate was set to 0.001.
@@ 3.2 Pixel CNN VAN and Gated PixelCNN VAN
In this section, we discuss the convolutional neural networks (CNN) version of the VAN [7]. It is based on the PixelCNN network used to generate images [16] and we will refer to it as PixelCNN VAN. The autoregressive property is assured by using a masked kernel (a kernel with some elements multiplied by zero).
@@
The original PixelCNN VAN [7] was designed for 2D systems, so we upgraded the code to 3D. In Fig. 2 we show a schematic view of the masked $3\times 3\times 3$ 3D kernel. The red, green, and blue colors denote different parts of the kernel, called stacks. The white elements are multiplied by 0 (masked) so that the autoregressive property is preserved. The black element is either zero (which means a strictly autoregressive layer) or non-zero (weakly autoregressive layer).
In reference [16] authors propose some improvements to the PixelCNN architecture. They note that the masking pattern leads to a blind spot in the receptive field even after applying many layers. To fix this, they propose to split the kernel into vertical and horizontal stacks. When adapting this architecture to 3D we have added the third depth stack (see Fig. 2). The information from depth is flowed into both vertical and horizontal stacks and from vertical to horizontal stack using 1$\times$1$\times$1 convolutions (see Figure 3 left). To ensure that the model is autoregressive, the depth stack does not receive information from the vertical and horizontal stacks, and the vertical stack is blinded to information from the horizontal stack. The output of each stack is then fed into the Exponential Linear Unit (ELU) activation function. Finally, a residual connection is added to the output of each stack, and a skip connection is drawn from the output of the horizontal stack. We follow [16] also in the use of the gated architecture: another mirror CNN produces a signal (number between zero and one) that is used to gate the output from the first CNN by multiplying it. Authors claim that this enhances the model performance. The architecture of the gated 3D PixelCNN is shown in the right part of Fig. 3. The input is a 3D spin configuration, as for VAN/HAN architectures. The network consists of several gated layers, where the first one is strictly autoregressive (s.a.), and the rest are weakly autoregressive. The final output is the sum of all the skip connections coming from layers fed into the sigmoid function to get the conditional probabilities of the spins.
The right panel of Figure 3 shows a schematic view of a single gated layer. To ensure that the model is autoregressive, the depth stack does not receive information from the vertical and horizontal stacks, and the vertical stack is blinded to information from the horizontal stack. The output of each stack is then fed into the Exponential Linear Unit (ELU) activation function. Finally, a residual connection is added to the output of each stack, and a skip connection is drawn from the output of the horizontal stack.
@@
We used the same hyperparameters for PixelCNN and gated PixelCNN: networks consist of 6 layers, with 18 channels and kernel $5\times 5\times 5$; an optimizer was ADAM and the learning rate was set to 0.003.
@@ 4.1 Timings and ESS for autoregressive architectures
Figure 4: Mean time (in seconds) of 10 batches generation of VAN (blue circles), HAN (orange circles), PixelCNN (red circles) and Gated PixelCNN (green circles) algorithms in dependence on system linear size $L$. Asymptotic approximations $\alpha L^{9}$ (blue dashed line), $\beta L^{6}$ (orange dashed line) and $\gamma L^{6}$ (green dashed line), for VAN, HAN and Gated PixelCNN accordingly, were superposed on plots. All the measurements were performed on Nvidia GeForce 4090 GPU.
@@
We shall start by comparing different sampling algorithms when it comes to the configuration generation time  [^4] Since in autoregressive networks, numerical cost of training is dominated by the generation of the configurations, the conclusions of this analysis apply to training timings as well.. We compare the generation time of 10 batches of $1024$ configurations for the architectures described in Section 3. In Fig. 4 we show the dependence on the linear system size $L$ on a log-log scale. First, we note that CNN-based architectures are significantly slower (an order of magnitude in the presented range of system sizes) than the VAN architecture based on dense networks. The HAN architecture is faster than VAN, as expected, and they have different scaling with system size, as we discussed in section 3.1. To better visualize the scaling of the architectures, we have added the lines in Fig. 4, which represent scaling $L^{9}$ (blue line) and $L^{6}$ (yellow and green line). For small system sizes, the scaling is modified by other effects (e.g. saving data), but for larger $L$ we observe the expected scaling for all algorithms. Also, the timing scaling of CNN architectures follows expected $\sim L^{6}$ behavior.
The final $ESS$ for Gated PixelCNN and HAN are similar, reaching around 0.95 for $L=8$ and close to the phase transition. For the PixelCNN this value is smaller, around 0.4. Finally, the VAN reaches around 0.6. Hence, we conclude that with current architectures, dense networks are more suitable for autoregressive networks. In what follows, we focus on the HAN algorithm as the most efficient.
@@ 4.2 Physical observables
We shall now cross-check the results obtained using HAN for some physical observable using MCMC. For this purpose, we use an in-house implementation of the Wolff cluster algorithm [3]. It is one of the state-of-the-art algorithms used in simulations of Ising model due to a significantly reduced (compared to local-update Monte Carlo like Metropolis algorithm) problem of critical slowing down close to the phase transition. At each step one builds a cluster of spins (where the probability of a spin to be attached to the cluster depends on $\beta$) and then the spins in the cluster are flipped.
@@
We start with the simulation of mean absolute magnetization per spin defined as:
$\left\langle{\left|M\right|}\right\rangle=\left\langle{\frac{1}{N}\left|\sum_{i=1}^{N}s^{i}\right|}\right\rangle_{p}.$ (14)
@@
In the limit of an infinite volume ($L\to\infty$), this quantity is an order parameter of the system with $\left\langle{\left|M\right|}\right\rangle=0$ for $\beta\to 0$ (disordered phase, high temperature), $\left\langle{\left|M\right|}\right\rangle=1$ for $\beta\to\infty$ (ordered phase, low temperature) and the phase transition at $\beta_{c}\approx 0.2215$.
@@
In Fig. 5 (left) we show $\left\langle{\left|M\right|}\right\rangle$ as a function of $\beta$ calculated for $L=8$ using two methods: HAN (with NIS) and Wolff algorithm. One can see very good agreement with the errors below 1‰, where the errors were obtained using the bootstrap method – see the inset which shows the difference of the two results. In the right panel of Figure 5 we show the mean energy $\left\langle{E}\right\rangle$, where a similar agreement can be observed.
@@
Figure 5: Comparison of the HAN (with NIS) with the Wolff algorithm for absolute magnetization per spin (left panel) and energy (right panel). Insets show difference of two results.
After cross-checking our results against the Markov chain Monte Carlo, we move to the observables that are not possible to calculate with this method. In Figure 6, we show results for the free energy per spin $F/N$ (8) and entropy per spin:
$S=\frac{\beta}{N}\left(\left\langle{E}\right\rangle-F\right),$ (15)
where both quantities were calculated with the NIS method. We show results for $L=4$ and $L=8$, demonstrating a weak dependence on the size of the system.
Figure 6: Free energy (left) and entropy (right) obtained using HAN (with NIS) for two system size: $L=4$ (blue) and $L=8$ (yellow).

### v2

@@ 3.1 Hierarchical autoregressive networks
The procedure for $L=16,32,64\ldots$ follows the same algorithm. Going from $k$ to $k+1$ (where $L=2^{k}$) requires: i) adding (on the second step) a new network and ii) increasing the size of the first network (first step), which fixes boundary spins. Snippets of the code that implement HAN 3D can be found in A.
Let us now calculate the total number of matrix multiplication operations that need to be performed to fix all the spins in the configuration using HAN. This would be asymptotically the leading computational cost of the algorithm. We assume 2 layers at each level of hierarchy, as we used in our implementation. For simplicity, we double-count some spins: for example, at the first level (red spins in Figure 1) we have $L^{2}+L(L-1)+(L-1)^{2}$ spins (3 walls), but we approximate this by $3L^{2}$ - this will obviously not affect the scaling at large $L$. As usual, $L$ is a power of $2$, but is not smaller than 4.
The first network in the hierarchy has $3L^{2}$ inputs and $3L^{2}$ outputs. We run the network $3L^{2}$ times, as this is the number of spins at this level of the hierarchy. Each spin generation requires double (we assumed two layers) multiplication of a $3L^{2}\times 3L^{2}$ matrix by a vector. So in total, we have $2\times 3L^{2}\times 3L^{2}$ multiplications per spin and $2(3L^{2})^{3}=54L^{6}$ operations in total.
At the second level of the hierarchy, one cross-like structure is fixed (green spins in Figure 1), containing $3L^{2}$ spins. At this level, the network has $3L^{2}+6L^{2}=9L^{2}$ inputs, where $6L^{2}$ is the number of spins surrounding the cross-like structure (in this case, it is the surface of the whole cube – note the redundancy here: we have $3L^{2}$ independent spins, but we feed them ”duplicated” to the network as $6L^{2}$ spins). The first layer has $9L^{2}$ inputs and $9L^{2}$ outputs, while the second layer has $9L^{2}$ inputs and $3L^{2}$ outputs. So, the number of operations at this level is $3L^{2}\times 9L^{2}\times 9L^{2}+3L^{2}\times 9L^{2}\times 3L^{2}=324L^{6}$.
At the third level of the hierarchy, one fixes 8 cross-like structures, each of size $3(L/2)^{2}$ spins. Each of the cross-like structures has $6(L/2)^{2}$ spins surrounding it. For a single cross-like structure, we can use the result from the previous paragraph, replacing $L\to L/2$: so we have $324(L/2)^{6}$ operations. For 8 cross-like structures, this would be $(324/8)L^{6}$ operations in total.
It is easy to see that the pattern will continue: going one step down the hierarchy, the number of operations gets a factor $1/8$ (the factor 8 due to more structures times the factor $1/64$ due to the smaller size of the cross-like structure). So, the number of multiplication operations from all the levels of the hierarchy for $L=2^{k},k\geq 2$ will be: $54L^{6}[1+6(1+8^{-1}+8^{-2}+8^{-3}+\ldots+8^{-(k-2)})]=54L^{6}\left[1+\frac{48}{7}(1-8L^{-3})\right]\approx(2970/7)L^{6}$, where in the approximation we skipped the sub-leading term since, from the beginning, we neglected them. [^5] The scaling with system size for the 3D hierarchical algorithm can also be determined using the master theorem for divide-and-conquer recurrences [19]. The problem has the following recurrence relation (for second and higher steps): $T(L)=8T(L/2)+\Theta(L^{6})$, which after applying the master theorem gives the scaling of $\Theta(L^{6})$. The first network is not a part of the recursion, but it also has a scaling $\Theta(L^{6})$, so overall the scaling remains $\Theta(L^{6})$.
For VAN, counting the number of multiplications is very simple: assuming two layers, we have two $L^{3}\times L^{3}$ matrices and the network needs to be evaluated $L^{3}$ times; so $2L^{9}$. Plugging in the value $L=8$ gives $2.7\times 10^{8}$ multiplications. This can be compared with the leading term for HAN, $(2970/7)8^{6}\approx 1.1\times 10^{8}$.
One should keep in mind that the scaling of $L^{6}$ in the number of multiplication operations does not automatically translate into the $L^{6}$ running time dependence, as other issues, e.g., parallelization on the graphics card, play an important role. Only in the limit of large $L$ we are expecting to see this behavior.
The HAN algorithm can also be implemented for sizes $L$ that are not powers of $2$. There are several possible extensions of this kind. Probably, the simplest one is to replace the heat bath algorithm (at the last stage of the hierarchy) by yet another network that generates a cube of spins (e.g., of size $2\times 2\times 2$). The same idea can be used to implement HAN for systems with cuboid shapes (non-cubic). We left those modifications for future work.
In the simplest implementation we present here, the unwanted (due to the autoregressive property) connections in the dense network are masked (weights multiplied by 0), but still, the operation of multiplication is performed. One may think of a different implementation, where, for a given conditional probability, only the necessary weights get multiplied. This trick can reduce the scaling of VAN and HAN: from $\sim L^{9}$ to $\sim L^{6}$ (for VAN) and from $\sim L^{6}$ to $\sim L^{4}$ (for HAN). However, the gains are not immediately obvious, as they depend on the size of the problem and the capacity of the GPU. Such an implementation is in progress and will be presented elsewhere.
Another neural-network-based approach, the TwoBo [5], has scaling $\sim L^{5}$, which is lower than HAN (in the current implementation). The RiGCS approach presented in Ref. [6] was shown to scale better than HAN in 2D; however, its version in 3D is not available. The comparison of those algorithms with the ones presented here is an interesting future direction of research. We should note that in HAN, due to sampling the boundary spins first, it is relatively easy to implement different boundary conditions. This is an important property for calculating observables such as classical mutual information or entanglement entropy (see Ref.[9, 10, 11]). To our knowledge, neither TwoBo nor RiGCS allows for such flexibility.
The results presented below were obtained with the same hyperparameters for the VAN and HAN networks: we always used dense networks with two layers: the first is followed by a PReLU activation function and the second by a sigmoid. The ADAM optimizer was chosen (see section 4.1 for discussion of learning rate choice). Inverse temperature $\beta$ was increased during training ($\beta$–annealing). We also used the $Z_{2}$ symmetry of the Ising model: $p(\mathbf{s})=p(-\mathbf{s})$: it allows for symmetrization of the probability distributions [7]: $q_{\theta}(\mathbf{s})\to[q_{\theta}(\mathbf{s})+q_{\theta}(-\mathbf{s})]/2\equiv\bar{q}(\mathbf{s})$. The batch size used to calculate the loss function and update the weights was 1000 configurations (the update of weights is called epoch). We stopped training when $ESS$ stabilized.
@@ 3.2 Pixel CNN VAN and Gated PixelCNN VAN
In this section, we discuss the convolutional neural networks (CNN) version of the VAN [7]. It is based on the PixelCNN network used to generate images [20]. The autoregressive property is ensured by using a masked kernel (a kernel with some elements multiplied by zero).
@@
The original PixelCNN VAN [7] was designed for 2D systems, so we upgraded the code to 3D. In Fig. 2 we show a schematic view of the masked $3\times 3\times 3$ kernel. The red, green, and blue colors denote different parts of the kernel, called stacks. The white elements are multiplied by 0 (masked) so that the autoregressive property is preserved. The black element is either zero (which means a strictly autoregressive (s.a.) layer) or non-zero (weakly autoregressive layer). In each autoregressive architecture, we use exactly one s.a. layer (the rest are weakly autoregressive) to ensure proper functional dependence of the conditional probabilities. In what follows, we shall call this architecture just PixelCNN since, except for autoregressive masking, it is based on standard 3D convolutional layers.
In Reference [20], the Authors propose some improvements to the PixelCNN architecture. They note that the masking pattern leads to a blind spot in the receptive field even after applying many layers. To fix this, they propose to split the kernel into vertical and horizontal stacks. When adapting this architecture to 3D, we have added the third depth stack (see Fig. 2). The information coming from the values of spins in the depth stack (denoted as $\rm{d}_{\rm{in}}$) flows into both the vertical ($\rm{v}_{\rm{in}}$) and the horizontal stacks ($\rm{h}_{\rm{in}}$) and from the vertical to the horizontal stack using $\times1$$\times$1 convolutions (see Figure 3 left). To ensure that the model is autoregressive, the depth stack does not receive information from the vertical and horizontal stacks, and the vertical stack is blinded to information from the horizontal stack. The output of each stack is then fed into the Exponential Linear Unit (ELU) activation function. Finally, a residual connection is added to the output of each stack ($\rm{d}_{\rm{out}}$, $\rm{v}_{\rm{out}}$ and $\rm{h}_{\rm{out}}$) and a skip connection is drawn from the output of the horizontal stack. In Ref. [20] the Authors propose also to use a gated architecture, which enhances the model performance. We shall follow this approach and call it gated Pixel CNN. The architecture of the gated 3D PixelCNN is shown in the right part of Fig. 3. The input is a 3D spin configuration, as for VAN/HAN architectures. The network consists of several gated layers, where the first one is strictly autoregressive (s.a.), and the rest are weakly autoregressive. The final output is the sum of all the skip connections coming from layers fed into the sigmoid function to get the conditional probabilities of the spins.
@@
We used the same hyperparameters for PixelCNN and gated PixelCNN: networks consist of 6 layers, with 18 channels and kernel $5\times 5\times 5$; the optimizer was ADAM and the learning rate was set to 0.003; batch size was 1000. We stopped the training when $ESS$ stabilized.
@@ 4.1 Timings and ESS for autoregressive architectures
Figure 4: Mean time (in seconds) of 10 batches generation using the VAN (blue circles), HAN (orange circles), PixelCNN (red circles) and Gated PixelCNN (green circles) algorithms in dependence on system linear size $L$. Asymptotic approximations $\sim L^{9}$ (blue dashed line), $\sim L^{6}$ (orange dashed line) and $\sim L^{6}$ (green dashed line), for VAN, HAN and Gated PixelCNN accordingly, were superposed on plots. All the measurements were performed on Nvidia GeForce 4090 GPU.
@@
We shall start by comparing different sampling algorithms when it comes to the configuration generation time  [^6] Since in autoregressive networks, numerical cost of training is dominated by the generation of the configurations, the conclusions of this analysis apply to training timings as well.. We compare the generation time of 10 batches of $1024$ configurations using the architectures described in Section 3. In Fig. 4 we show the dependence on the linear system size $L$ on a log-log scale. First, we note that CNN-based architectures are significantly slower (an order of magnitude in the presented range of system sizes) than the VAN architecture based on dense networks. The HAN architecture is faster than VAN, as expected, and they have different scaling with system size, as we discussed in section 3.1. To better visualize the scaling of the architectures, we have added the lines in Fig. 4, which represent scaling $\sim L^{9}$ (blue line) and $\sim L^{6}$ (yellow and green line). For small system sizes, the scaling is modified by other effects (e.g. saving data or parallelization), but for larger $L$ we observe the expected scaling for all algorithms except HAN. Also, the timing scaling of CNN architectures follows the expected $\sim L^{6}$ behavior.
Figure 5: Moving average of ESSESS for different models as a function of epoch (left panel) and time (right panel) for $\beta=0.22$, $L=8$. Curves with the same colors show different runs.
To compare the training efficiency between different methods, we present in Figure 5 how $ESS$ changes during their training for $L=8$ and $\beta=0.22$ (which is close to the critical temperature for the 3D Ising model – see next subsection). We perform measurements of the $ESS$ at each batch of configurations that were used to calculate the loss function. Since the batch is small ($\sim 1000$ samples), the statistical fluctuations of $ESS$ can be large; therefore, we perform a simple moving average over 50 epochs to get smoother curves. We ran the algorithms several times to check if the results are stable (group of curves with the same colors). For the purpose of this comparison, we used one learning rate (lr) for VAN and HAN, lr=0.0003. Note that usually, to get the best performance, one tunes the lr for each architecture separately. However, then the comparison of how $ESS$ changes with epochs is difficult to interpret. In the Pixel CNN and the gated Pixel CNN networks, we used lr=0.003, which was checked to be optimal for them - using the same lr as for dense networks would significantly slow down the training. On the other hand, for dense architectures, lr=0.003 is too large and the final $ESS$ is smaller than for lr=0.0003. The total number of epochs for each architecture was chosen such that the $ESS$ reaches its maximum.
In the left panel of Figure 5 we show the $ESS$ as a function of the number of epochs. We see that CNN architectures (pixel CNN and gated pixel CNN) need fewer epochs to be trained than dense architectures (VAN and HAN). This is mostly due to the fact that lr for those architectures can be 10 times larger. One can also see that HAN needs $\sim 2$ times fewer epochs than VAN and reaches a better final $ESS$. The situation is quite different if we plot the $ESS$ as a function of the time of training (right panel of Figure 5 ). Since the generation of configurations in the gated Pixel CNN is much slower than in HAN, the HAN algorithm reaches its maximal $ESS$ in a much shorter time. We also notice that for Pixel CNN, the curve of training depends very much on the run; nevertheless, the final $ESS$ is always much lower than the values obtained with other models.
System size VAN HAN PixelCNN Gated PixelCNN
4 95s 26s 84s 38s
6 794s - not reached 991s
8 4689s 1078s not reached 18668s
Table 1: Time of training needed for moving average ESS to reach $0.7$ for $\beta=0.22$, see the text for more details. The values were shown for one run, they can vary up to 10% for different runs.
We shall now demonstrate how the training time grows with the system size. For this purpose, we measure the time of a given architecture’s training to the value of ESS=0.7ESS=0.7.777The statistical error of the observables calculated using NIS is proportional to $1/\sqrt{ESS}$, therefore, the value of 0.7 leads to $20\%$ larger error than for a perfect sampler ($q_{\theta}=p$) – for this comparison we assume it is a sufficient level to say that the network is well-trained. In Table 1 we show the timings obtained for $\beta=0.22$ and for three system sizes: $L=4,6,8$. Since the moving average of ESS has some statistical fluctuation, the value presented is the mean between the first time the threshold value 0.7 was reached by the moving average and the last time when the $ESS$ was below that value. The learning rates were the same for all system sizes and were chosen as described in the paragraph above (lr=0.0003 for VAN/HAN and lr=0.003 for PixelCNN/Gated Pixel CNN).
First, we observe that for all architectures the time of training grows much faster than the time of generation of batch presented in Figure 4 (note that in the range $L\in[4,8]$ VAN and HAN do not follow their asymptotic scalings yet). This is just due to the fact that when increasing the size of the system, the networks need more epochs to be trained. As was shown for normalizing flow architectures in 2D, the numerical cost of training may scale exponentially with system size [21]. Here, we cannot confirm such behavior as larger system sizes in 3D are currently out of reach due to small ESS values obtained.
For VAN, we observe a factor of $\sim 50$ increase in training time by increasing the linear size of the system by a factor of 2; for HAN, this factor is $\sim 40$. The fastest growth of training time is observed for Gated Pixel CNN, where the timing grows by a factor of $\sim 500$ between $L=4$ and $L=8$. Finally, the standard Pixel CNN architecture did not reach ESS=0.7 when applied to $L\geq 6$.
To check the maximal $ESS$ the networks can reach at $\beta=0.22$, we tune the lr for each architecture separately. For VAN and HAN, the lr was changed during the training from 0.005 to 0.0001 manually (we decreased lr when no progress in $ESS$ was observed). Finally, the VAN reaches $ESS$ of 0.8 and HAN of 0.95. (measured at large statistics). For Gated PixelCNN, $ESS$ is similar to that for HAN, reaching around 0.95. For the PixelCNN, this value is smaller, around 0.4.
Using the same statistics, the difference in statistical uncertainties between different methods is not large (for example, the error for PixelCNN is $\sqrt{0.95/0.4}\approx 1.5$ larger than for HAN). In fact, the main difference between the methods is in the generation time of a given statistic (presented in Fig. 4). In what follows, we focus on the HAN algorithm as it reaches very good $ESS$ and is faster than other methods. However, we confirmed that all the methods give the same results for physical observables, within the uncertainties.
@@ 4.2 Physical observables
We shall now cross-check the results obtained using HAN for some physical observables using MCMC. For this purpose, we use an in-house implementation of the Wolff cluster algorithm [3]. It is one of the state-of-the-art algorithms used in simulations of the Ising model due to a significantly reduced (compared to local-update Monte Carlo like the Metropolis algorithm) problem of critical slowing down close to the phase transition. At each step, one builds a cluster of spins (where the probability of a spin to be attached to the cluster depends on $\beta$) and then the spins in the cluster are flipped. The HAN was trained to $ESS=0.95$ and then $20\times 10^{6}$ configurations were generated for each inverse temperature $\beta$. For the Markov Chain generated using the Wolff algorithm, $20\times 10^{6}$ configurations were obtained to calculate observables.
@@
We start with calculating the mean absolute magnetization per spin, defined as:
$\frac{\left\langle{\left|M\right|}\right\rangle}{N}=\left\langle{\frac{1}{N}\left|\sum_{i=1}^{N}s^{i}\right|}\right\rangle_{p}.$ (14)
@@
In the limit of an infinite volume ($L\to\infty$), this quantity is an order parameter of the system with $\left\langle{\left|M\right|}\right\rangle/N=0$ for $\beta\to 0$ (disordered phase, high temperature), $\left\langle{\left|M\right|}\right\rangle/N=1$ for $\beta\to\infty$ (ordered phase, low temperature) and the phase transition occurs at $\beta_{c}\approx 0.2215$.
@@
In Fig. 6 (left), we show $\left\langle{\left|M\right|}\right\rangle/N$ as a function of $\beta$ calculated for $L=8$ using two methods: the HAN (with NIS) and the Wolff algorithm. One can see very good agreement with the errors below 1‰, where the errors were obtained using the bootstrap method – see the inset which shows the difference of the two results. In the right panel of Figure 6, we show the mean energy per spin $\left\langle{E}\right\rangle/N$, where a similar agreement can be observed.
@@
Figure 6: Comparison of the HAN (with NIS) with the Wolff algorithm for absolute magnetization per spin (left panel) and energy per spin (right panel). Insets show the difference between the two results; its uncertainty was obtained by adding the Wolff and HAN errors in quadrature.
We now move to the thermodynamical observables. In Figure 7, we show results for the free energy per spin $F/N$ (8) and entropy per spin:
$\frac{S}{N}=\frac{\beta}{N}\left(\left\langle{E}\right\rangle-F\right),$ (15)
where both quantities were calculated using the NIS method. We show results for $L=4$ (red dashed curve) and $L=8$ (orange points), demonstrating a weak dependence on the size of the system. To cross-check our results, we implemented the Wang-Landau algorithm (W-L) [22], which is a random walk algorithm capable of estimating the density of states. In Figure 7, we show the results of W-L for L=8 with a solid blue line. The agreement between HAN and W-L is very good (see insets).
The results for W-L were obtained by averaging 1000 independent runs. In each run, the parameter of W-L $\ln f$ was halved when the flatness of the histogram reached 95%. The runs were started with $\ln f=1$ and stopped when $\ln f<10^{-9}$. We used a dual-socket AMD EPYC 7F52 CPU (64 hardware threads in total) and the data presented in Figure 7 were calculated within $\sim 10^{4}$s.
The HAN $L=8$ results from Figure 7 were obtained after $\sim 1000$s of training and $\sim 100$s of sample generation at each $\beta$ value; a GeForce RTX 4090 GPU was used. At $\beta=0.22$, the W-L reaches a similar statistical error (as HAN) for $S/N$ in $\sim 3000s$. However, to get similar errors for $F/N$, one would need to run the W-L for $\sim 10^{6}$s because the uncertainty of HAN for $F/N$ is much lower than for $S/N$. Although HAN is significantly faster, one should note that this is a very rough comparison; e.g., due to different hardware used to run both algorithms and a possible choice of a more optimal schedule for W-L. Also, in W-L, once the density of states is obtained, calculating the results for all temperatures is instantaneous, whereas HAN requires training at each temperature.
Figure 7: Free energy per spin (left) and entropy per spin (right) as a function of inverse temperature $\beta$. Results obtained using HAN (with NIS) were shown for two system sizes: $L=4$ (red dashed curve) and $L=8$ (orange points). For $L=8$, results from the Wang-Landau algorithm [22] were presented (blue solid curve). Insets show the difference between the two results; its uncertainty was obtained by adding the W-L and HAN errors in quadrature.
As stated above, NIS requires that the supports of $q_{\theta}$} and $p$ coincide. If they differ, we are talking about mode collapse, which is most often caused by broken symmetries (see [23] for the discussion). Mode collapse can occur in reverse training, since the network generates its own training data (spin configurations). When a mode is missing in $q_{\theta}$, the $ESS$ can still be close to 1 because the configurations where $q_{\theta}$ seriously underestimates $p$ will never appear in the sampling.
We did not observe this mode-seeking behavior for the networks we trained; this is due to the $Z_{2}$ symmetry used in the training (see Section 3.1). In the absence of mode collapse, the NIS estimator of $Z$ given by equation (11) is unbiased. However, it has a non-zero variance ($ESS\neq 1$) due to imperfect training of the network. Because of this variance, estimators of other quantities that are functions of $Z$, like the free energy $F$ or entropy $S$, will usually be biased - see the Supplementary Materials of Ref. [24] or the Appendix of Ref. [9]. However, as was shown in Ref. [24], the bias decreases with statistics as $1/\cal N$; hence, for large statistics (and good $ESS$) the bias is negligible compared to the statistical error ($\propto 1/\sqrt{{\cal N}}$). To conclude, the systematic errors of NIS in the simulations we performed are insignificant.